2)   A company makes an investment in a new delivery van costing $47,000 and is expected to have a service life of 200,000 miles at an annual usage of 40,00 miles per year.  The delivery service is expect to generate the following profits:  

$15K, in odd years and,  - $20K, in even years, and the project is evaluated at a rate of 6%,

  1. What is the present value of this investment?
  2. What is the annual equivalent?

n          In                     Out                                                                

            0                                  30000                                                             

            1          40000             15000                                                             

            2          40000             15000                                                             

            3          40000             15000                                                             

            4          40000             15000                                                             

            5          40000+3000                                                                                      
What is the Present value of the cash flow above
Answer #1

Answer 2 A)

Present worth of Investment =PW of Profits - Initial costs

=(15000/1.06)+(20000/1.06^2)+(15000/1.06^3)+(20000/1.06^4)+(15000/1.06^5)-47000

=71595.90 - 47000

=24595.9

Answer 2 B)

Annual equivalent=Present worth/Sum of discounting factors = 24595.9 /4.2125 = $5838.98

Note: for the last question, rate is required npv cannot be calculated without rate or cost of capital.